A ball thrown with a speed of 100m/s attains a height of 150m calculate the time of flight the angle of projection the range(10m/s^2
1
Expert's answer
2020-11-05T10:38:37-0500
"2gh=v^2\sin^2{\theta}\\2(10)(150)=100^2\sin^2{\theta}\\\theta=33\degree"

"t=\frac{2(100)\sin{33}}{10}=11\ s"

"R=\frac{100^2\sin{66}}{10}=914\ m"